Circular No. 8249

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2P/ENCKE F. Bensch, Harvard Smithsonian Center for Astrophysics (CfA); E.

Bergin, University of Michigan; and G. Melnick, CfA, write: “We have monitored the 1(10)-1(01) emission of water vapor at 556.936 GHz toward comet 2P using the Submillimeter Wave Astronomy Satellite (SWAS). Between Nov. 7.04 and 15.21 UT, the line-integrated antenna temperature within the 3′.2 × 4′.5 elliptical SWAS beam varied between I < 0.252 (3σ upper limit for observations on Nov. 7.04–7.99) and I = 0.98 K km/s (Nov. 12.03–12.99). The average line-integrated intensity for this period is I = 0.55 ± 0.03 K km/s. The water-production rate, Q(H_{2}O), is derived using a spherical outflow model (Haser model) with a water photo-destruction rate of 1.366 × 10^{−5} s^{−1} and an assumed ortho-para ratio of 3. The uncertainty in the resulting Q(H_{2}O) is governed by the finite S/N ratio of the observations and the uncertainty in the electron abundance in the coma. (In addition to infrared fluorescence and H_{2}O-H_{2}O collisions, H_{2}O-electron collisions provide a significant contribution to the 1(10)-1(01) line excitation; our modeling of the electron abundance uses the same parameterization as Biver et al. 1999, A.J. 118, 1850). For observations made between Nov. 9.06 and 9.96 (I = 0.53 ± 0.07 K km/s), we derive Q(H_{2}O) = (2.9 ± 0.4) × 10^{27} s^{−1} for an electron abundance similar to those derived by in-situ measurements in the coma of 1P/Halley, and Q(H_{2}O) = (4.0 ± 0.5) × 10^{27} s^{−1} for an electron abundance reduced by a factor of 0.2. Previous studies of this transition toward several other comets by SWAS and by the (sub)millimeter-wavelength satellite Odin have indicated that the electron density in cometary comae might be smaller by a factor of ∼ 0.2, compared to the electron density in 1P/Halley (Biver, private communication, based on data from Lecacheux et al. 2003, A.Ap. 402, L55).”

Recent visual total-magnitude and coma-diameter estimates: Nov. 17.11 UT, 7.3, 22′ (C. S. Morris, Fillmore, CA, 10×50 binoculars); 18.83, 6.9, 25′ (E. P. Bus, Moncarapacho, Portugal, 10×56 binoculars); 21.76, 7.3, 12′ (P. Guzik, Krosno, Poland, 10×50 binoculars); 24.10, 7.2, 15′ (Morris); 26.72, 6.8, 11′ (A. Kammerer, Ettlingen, Germany, 9×63 binoculars); Dec. 3.08, 6.7, 8′ (Morris, 20×80 binoculars); 3.96, 6.2, ∼ 8′ (J. E. Bortle, Stormville, NY, 15×70 binoculars).

NOVAE IN M31 Corrigendum. On IAUC 8249, line 9, for 29.883, 19.9. read 29.883,

18.9.

2003 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green